OpenStudy (anonymous):

Hi can someone help me with this ? Write a vector equation and parametric equations of lines: The line is the intersection of the planes x + y + z = 1 and 3x-7y + z = 11

ganeshie8 (ganeshie8):

you need to find the equation of the intersection of given planes

ganeshie8 (ganeshie8):

for that you need below : 1) direction cosines 2) any point on the line

OpenStudy (anonymous):

I started off by minusing the 2 equations

ganeshie8 (ganeshie8):

you can find direction cosines by the fact that the "line" is perpendicular to both the "planes normals"

ganeshie8 (ganeshie8):

start by finding direction cosines

OpenStudy (anonymous):

I ended up getting x = 5+4y and z = -4 -5y ..

OpenStudy (anonymous):

the answer is r=5i-4k + t(4i+j-5k).. I got the \[r _{0}\] but i don't know how to get the t(4i+j-5k)

ganeshie8 (ganeshie8):

start by finding direction cosines

OpenStudy (anonymous):

I'm confused ..

ganeshie8 (ganeshie8):

its easy, lets start from beginning :)

ganeshie8 (ganeshie8):

first step : observe that you're given two "planes"

ganeshie8 (ganeshie8):

and you're asked to find the equation for intersection of those planes.

ganeshie8 (ganeshie8):

thats the problem at hand, ok ?

OpenStudy (anonymous):

Alright

ganeshie8 (ganeshie8):

wat do u knw about equation of plane ?

ganeshie8 (ganeshie8):

in the plane equation : \(ax + by + cz = d\) the direction consines of its normal are : \(\left< a, b, c\right >\)

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

Agreed

ganeshie8 (ganeshie8):

and also we knw that, EVERY line on the plane is "perpendicular" to the plane's normal

OpenStudy (anonymous):

Right

ganeshie8 (ganeshie8):

good, next see that the required line lies on both the planes

ganeshie8 (ganeshie8):

so, it is "perpendicular" to both the plane's normals

ganeshie8 (ganeshie8):

so, if u take "cross product" of given planes' normals, you wud get the direction cosines of the required line

ganeshie8 (ganeshie8):

planes : x + y + z = 1 and 3x-7y + z = 11

OpenStudy (anonymous):

But how do you do that when you have a d ?

ganeshie8 (ganeshie8):

direction cosines of any line perpendicular to "normals of these planes" = <1, 1, 1> x <3, -7, 1>

ganeshie8 (ganeshie8):

d will not affect the direction. d oly changes the distance from origin.

OpenStudy (anonymous):

so we get 8 2 -10?

OpenStudy (anonymous):

Then when do we integrate d in our calculations ?

ganeshie8 (ganeshie8):

yes, thats the direction vector of the required line.

ganeshie8 (ganeshie8):

you just need some "point on the line", then u can write the equation of line as : r = <point> + t<8, 2, -10>

OpenStudy (anonymous):

Which i already found at the start .. awesome!

OpenStudy (anonymous):

I substracted both equations, got rid of z, replaced it in the first equation and got z = -4-5y .. and x = 5 + 4 y ... I think I answered to fast lol I just confused myslef i think

ganeshie8 (ganeshie8):

another easy way for finding the "some point" is to simply plugin z = 0 u wud get two equations in x and y solve them

ganeshie8 (ganeshie8):

then, the "some point" wud be : <x1, y1, 0>

ganeshie8 (ganeshie8):

u may use which ever method u feel comfortable... :)

OpenStudy (anonymous):

i got x = 5 and y = -4

OpenStudy (anonymous):

ButWhich gives d(5 , -4 , 0) but my answer says its's ( 5, 0, -4)

ganeshie8 (ganeshie8):

plugin ur point in given plane equations, does it satisfy both equations ?

OpenStudy (anonymous):

Did i miss something ? lol

OpenStudy (anonymous):

Ohh i got it cool thanks again :D

ganeshie8 (ganeshie8):

u wlc :)